Question

For the influenza season 2014-15, there was some question about the effectiveness of the vaccine. In order to determine whether the vaccine even had an effect in preventing the flue, we gathered the following:

Did you get the flu in 2014-15?

Were you Vaccinated

Yes

No

Totals

No

332

580

912

Yes

621

1254

1875

Totals

953

1834

2787

Ok…. So with this data:

  1. What is the probability that you got the flu given you were not vaccinated? I.e. the Probability (got the flu │ you were not vaccinated)? (1)
  2. What is the probability that you got the flu given you were vaccinated? I.e. the Probability (got the flu │ you were vaccinated)? (1)
  3. What is the Relative Risk? (2)
  4. Interpret (1)
Answer #1

Please don't hesitate to give a "thumbs up" for the answer in case the answer has helped you

a. P(given not vaccinated, got flu) = had flu when not vaccinated/ total not vaccinated = 332/912 = .3640

b. P(given vaccinated , got flu) = had vaccinated and got flu / total vaccinated = 621/1875 = .3312

c. relative risk = (621/1875) / (332/912) = .3312/.3640 = .9099